# Adaptive FEM with explicit time integration for the wave equation

### v1

@@ 1 Introduction
A posteriori error estimates are the cornerstone of any adaptive strategy with guaranteed rigorous error bounds. For elliptic problems, a posteriori error analysis is well-developed and leads to rigorous and explicitly computable error bounds, which allow the end user to assess the accuracy of $u_{h}$, see [AinsworthOden:00:book:A-posteriori, Verfurth:13:book:A-posteriori] and references therein. Typically, such error bounds consist of local contributions, known as error indicators, which can be used to automatically steer local mesh refinement strategy and thus improve upon the accuracy of $u_{h}$ while keeping the added cost minimal. For parabolic problems, both time-discretization and mesh change must also be included in the a posteriori error estimates. Various a posteriori error bounds are available for parabolic problems either based on duality or space-time Galerkin formulation, often coupled with a discontinuous Galerkin (DG) formulation in time [ErikssonJohnson:91:article:Adaptive, Picasso:98:article:Adaptive, Verfurth:03:article:A-posteriori, ChenJia:04:article:An-adaptive, Verfurth:13:book:A-posteriori, GaspozSiebertKreuzerZiegler:19:article:A-convergent].
@@
For (time-dependent) wave equations, error estimation and adaptivity are far less developed than for elliptic or parabolic problems. Indeed, in contrast to parabolic problems, the most commonly used time integration methods for wave equations, such as the popular second-order leapfrog (LF) method (or Störmer-Verlet), are explicit. Explicit time integration is highly efficient, in particular for large-scale problems on massively parallel architectures. In the presence of local mesh refinement, however, the CFL stability constraint will impose a tiny time-step across the entire computational domain which will cripple any explicit time integrator; hence, standard explicit methods are generally deemed inefficient when combined with adaptivity as ”for adaptive methods to time-dependent waves, unconditionally stable time-stepping methods are needed.” ([ThompsonHe:05:article], p. 1948)
@@
It is no wonder thus that most a posteriori error estimates and space-time adaptive strategies for the wave equation have relied on implicit time integration [Johnson:93:article:Discontinuous, HulbertHughes:90:article:Space-Time-Finite, BernardiSuli:05:article:Time]. Residual based a posteriori error estimates with first-order implicit time-stepping were developed in [BernardiSuli:05:article:Time, Adjerid:02:article:A-posteriori, Adjerid:06:article:A-posteriori]. Alternatively, goal-oriented adaptivity always requires the solution of an adjoint (dual) problem [BangerthRannacher:01:article:Adaptive, BangerthGroteHohenegger:04:article]. Recently, various a posteriori error estimates were derived for semi-discrete formulations with anisotropic mesh refinement either in the $L^{2}(0,T;H^{1}(\Omega)$-norm [Picasso:10:article:Numerical, GoryninaLozinskiPicasso:19:article:An-Easily-Computable] or in a “damped energy norm” [Chaumont-Frelet:23:article:Asymptotically-Constant-Free]; the latter recently led to fully discrete error estimates [Chaumont-FreletErn:24:techreport:Damped-Energy-Norm] yet under a fixed mesh assumption. An unconditionally stable space-time Galerkin discretization of the second-order wave equation was also derived in [ferrari2025intrinsicunconditionalstabilityspacetime].
@@
In the design of space-time adaptive methods, a posteriori error estimates of fully discrete formulations in time-dependent problems need to include the added effect on the error due to mesh change from one time-step to the next; in fact, some of the above cited works address that particular issue for parabolic problems – see also [Dupont:82:article:Mesh, LakkisPryer:12:article:Gradient, SchmidtSiebert:05:book:Design]. Both for accuracy and efficiency, it is indeed expected, often even required, for an adaptive method to locally refine and coarsen the mesh repeatedly during the entire simulation. Although quantification of mesh-change error in second order hyperbolic problems is less studied, a notable result in this direction was provided by [KarakashianMakridakis:05:article:Convergence] in an a priori setting.
@@
To overcome the above mentioned difficulties, we recently derived a posteriori error estimates for the wave equation with explicit LF time integration and mesh change [GroteLakkisSantos:24:PosterioriErrorEstimatesWave], starting from the a posteriori analysis for semi-discrete formulations [GeorgoulisLakkisMakridakisVirtanen:16:article:A-Posteriori] (continous in space) discretized in time using a two-step Newmark (or cosine) family of methods, which include the explicit leapfrog method. Moreover, our estimates also allow for LF based explicit local time-stepping (LF-LTS), which circumvents the local CFL bottleneck by taking smaller local time-steps, but only where needed, without sacrificing explicitness.
@@
Here we propose an adaptive strategy for the solution of the time-dependent wave equation based on rigorous a posteriori error estimates which include both mesh change and LF based fully explicit local time-stepping. In Section 2, we present the finite element discretization including the Galerkin FE formulation with explicit LF time integration and mesh change. In Section 3, we develop an adaptive method for the numerical solution of the wave equation with combines three essential ingredients: a posteriori estimates for guaranteed error bounds, mesh change to adapt the mesh from one time-step to another, and local time-stepping to retain the efficiency of explicit time integration even in the presence of local mesh refinement. In fact, local error indicators not only permit to estimate but also to control the effect of mesh change and adaptivity and thus to decide ”on the fly” whether to accept or recompute the numerical solution at the current time-step. By restricting mesh change to compatible meshes, as detailed in Section 3.2, we ensure that local refinement never increases the numerical error while keeping under control the inherent but small error from mesh coarsening. Finally, in Section 4, we present three numerical experiments which confirm the expected convergence order and illustrate the usefulness of our adaptive strategy.
@@ 2 Galerkin FE formulation with explicit time integration and mesh change
For $u_{0}\in H^{1}_{\Gamma_{0}}(\Omega)$, $v_{0}\in L^{2}(\Omega)$ and $f\in L^{2}(0,T;L^{2}(\Omega))$, it is well-known that the wave equation (2.1) in fact has a unique solution $u\in C^{0}(0,T;H^{1}_{\Gamma_{0}}(\Omega))\cap C^{1}(0,T;L^{2}(\Omega))$ [LionsMagenes:72:.Non-HomogeneousBoundaryValueProblemsandApplications].
@@
$h^{n}(x):=\operatorname{diam}\bigcap_{\boldsymbol{x}\in K\in\mathcal{M}_{n}}\bar{K}$ (2.6)
and we denote by $h_{E}$ the constant value of $h^{n}$ on each $E\in\mathcal{M}_{n}\cup\operatorname{Sides}\mathcal{M}_{n}$.
With adaptvity in mind, we allow for the case where $\mathcal{M}_{n}$ (and thus $\mathbbmss{V}_{n}$) changes with time under mesh compatibility conditions, which implies that at each point of the domain either $\mathcal{M}_{n-1}$ is a (compatible) refinement of $\mathcal{M}_{n}$ or conversely – see Section 3.2 and [LakkisMakridakis:06:article:Elliptic] for details.
@@
To transfer functions between FE spaces during mesh change, for instance, we introduce the FE transfer operator $\Pi_{n}:\mathrm{C}^{0}(\Omega)\rightarrow\mathbbmss{V}_{n}$, which may either denote the $L^{2}$-projection or standard Lagrangian interpolation
@@
thanks to Riesz’s representation theorem by
@@
The particular instance of $\widetilde{A_{n}}$ in (2.16) corresponds to the simplest situation with two local time-steps of size $\Delta t/2$ each for each global time-step of size $\Delta t$. By letting $\widetilde{A_{n}}$ denote a generic perturbed bilinear form induced by local time-stepping, our analysis inherently encompasses situations with different coarse-to-fine time-step ratios, too, which may even change from one locally refined subregion to another across a single mesh. In fact, it even includes a hierarchy of locally refined regions, each associated with its own local time-step [DiazGrote:15:article:Multi-Level-Explicit]
@@
$U^{0}$ $:=\Pi_{0}u_{0}$ (2.17)
$U^{1}$ $:=\Pi_{1}\left[U^{0}+\Pi_{0}v_{0}\,\Delta t+\left(F^{0}-\widetilde{A_{0}}U^{0}\right)\Delta t^{2}\right]$
$U^{n+1}$ $:=\Pi_{n+1}\left[2\,U^{n}-\Pi_{n}U^{n-1}+\left(F^{n}-\widetilde{A_{n}}U^{n}\right)\Delta t^{2}\right],\>n=1,\ldots,N.$
@@ 3 Space-time adaptive finite element method
We shall now develop a fully adaptive method for the numerical solution of the wave equation with combines three essential indredients: a posteriori estimates for guaranteed error bounds, mesh change to adapt ”on the fly” the mesh from one time-step to another, and local time-stepping to retain the efficiency of explicit time integration even in the presence of local mesh refinement.
@@ 3.1 A posteriori error estimates
In [GroteLakkisSantos:24:PosterioriErrorEstimatesWave] we derived the following a posteriori error estimates for the discretized time-dependent wave equation (2.17) with mesh change:
@@
$\max_{1\leq n\leq N}\|V^{n-\frac{1}{2}}-v(.,t_{n-\frac{1}{2}})\|_{\operatorname{L_{2}}(\Omega)}\\ \leq\|\boldsymbol{e}(0)\|_{\text{erg},\mathcal{A}}+C\left\{2\sum_{m=1}^{2N}\zeta^{m}+\max_{1\leq n\leq N}\varepsilon_{1}^{n}\right\}.$ (3.2)
Here the potential energy norm is given by
@@
$\langle\phi,\chi\rangle_{\mathrm{erg},\mathcal{A}}:=\langle\mathcal{A}\phi_{0}\mid\chi_{0}\rangle+\langle\phi_{1},\chi_{1}\rangle,\qquad\phi=\begin{bmatrix}\phi_{0}\\ \phi_{1}\end{bmatrix},\,\chi=\begin{bmatrix}\chi_{0}\\ \chi_{1}\end{bmatrix}\in H^{1}_{\Gamma_{0}}\times\operatorname{L_{2}}(\Omega).$ (3.5)
The upper bounds in (3.1) and (3.2) involve standard a posteriori error estimators for elliptic problems with respect to the energy or the $L^{2}$- norm [Verfurth:13:book:A-posteriori, AinsworthOden:00:book:A-posteriori] together with additional terms due to time integration or mesh change:
@@
$\vartheta_{1}^{n}(t)$ $:=\Delta t^{2}\begin{cases}\left\|\frac{1}{2}\partial^{2}U^{n}\ell_{n}(t)-\partial V^{n-\frac{1}{2}}q_{n-\frac{1}{2}}(t)\right\|_{\operatorname{L_{2}}(\Omega)},\qquad t\in I^{\prime}_{n},\\[10.00002pt] \left\|\frac{1}{2}\partial^{2}U^{n}\ell_{n}(t)-\partial V^{n-\frac{1}{2}}q_{n+\frac{1}{2}}(t)\right\|_{\operatorname{L_{2}}(\Omega)},\qquad t\in I^{\prime}_{n+\frac{1}{2}},\end{cases}$ (3.13)
@@ 3.2 Compatible mesh change
Mesh change during any time-dependent simulation cannot be arbitrary without risking a significant loss in accuracy [Dupont:82:article:Mesh]. Indeed both mesh coarsening or refinement will generally increase the numerical error when transferring the FE solution via interpolation or projection from one mesh to another. To ensure that local mesh refinement never leads to a loss in accuracy, we restrict mesh change to compatible meshes only. Moreover, we shall allow coarsening only where appropriate to minimize the resulting inherent information loss.
@@
$\beta_{K}^{0}$ ::=:= $\left\|\left[\Pi_{n+1}-\operatorname{Id}\right]y\right\|_{\mathcal{A},K}=\left\|(y-\hat{y})|_{K}\right\|_{\mathcal{A}}$ (3.18)
$\beta_{K}^{1}$ ::=:= $\left\|\left[\Pi_{n+1}-\operatorname{Id}\right]y\right\|_{L^{2}(K)}=\left\|(y-\hat{y})|_{K}\right\|_{L^{2}}.$ (3.19)
@@
Following [LakkisPryer:12:article:Gradient] we now exemplify this point by considering the simplest situation of a one-dimensional FE approximation with continuous piecewise linear polynomials. Again let $K_{-}=(x_{m-1},x_{m})$, $K_{+}=(x_{m},x_{m+1})$ both marked for coarsening and thus possibly replaced by the single element $\widehat{K}=(x_{m-1},x_{m+1})$. For $x\in\widehat{K}$ we have
@@
$y(x)-\hat{y}(x)=\left[y_{n}-\frac{1}{2}(y_{n-1}+y_{n+1})\right]\Phi_{m}^{n}(x),\qquad x\in\widehat{K},$ (3.20)
@@
$\beta_{K_{-}}^{1}$ $=$ $\left|y_{n}-\frac{1}{2}(y_{n-1}+y_{n+1})\right|\left\|\Phi^{n}_{m}\right\|_{L^{2}{(K_{-})}}$
$\beta_{K_{+}}^{1}$ $=$ $\left|y_{n}-\frac{1}{2}(y_{n-1}+y_{n+1})\right|\left\|\Phi^{n}_{m}\right\|_{L^{2}{(K_{+})}},$
@@
By allowing for coarsening only in those elements where the pre-indicators are small, we thus also avoid any detrimental drastic increase in the mesh-change indicators $\mu_{0}^{n},\mu_{1}^{n}$ and $\mu_{2}^{n}$ from Section 3.1 .

### v3

@@ 1 Introduction
A posteriori error estimates are the cornerstone of any adaptive strategy with guaranteed rigorous error bounds. For elliptic problems, a posteriori error analysis is well-developed and leads to rigorous and explicitly computable error bounds, which allow the end user to assess the accuracy of $u_{h}$, see [undefb, undefan] and references therein. Typically, such error bounds consist of local contributions, known as error indicators, which can be used to automatically steer local mesh refinement strategy and thus improve upon the accuracy of $u_{h}$ while keeping the added cost minimal. For parabolic problems, both time-discretization and mesh change must also be included in the a posteriori error estimates. Various a posteriori error bounds are available for parabolic problems either based on duality or space-time Galerkin formulation, often coupled with a discontinuous Galerkin (DG) formulation in time [undefq, undefag, undefam, undefl, undefan, undefs].
@@
For (time-dependent) wave equations, error estimation and adaptivity are far less developed than for elliptic or parabolic problems. Indeed, in contrast to parabolic problems, the most commonly used time integration methods for wave equations, such as the popular second-order leapfrog (LF) method (or Störmer-Verlet), are explicit. Explicit time integration is highly efficient, in particular for large-scale problems on massively parallel architectures. In the presence of local mesh refinement, however, the CFL stability constraint will impose a tiny time-step across the entire computational domain which will cripple any explicit time integrator; hence, standard explicit methods are generally deemed inefficient when combined with adaptivity as ”for adaptive methods to time-dependent waves, unconditionally stable time-stepping methods are needed” ([undefal], p. 1948).
@@
It is no wonder thus that most a posteriori error estimates and space-time adaptive strategies for the wave equation have relied on implicit time integration [undefab, undefaa, undefe]. Residual based a posteriori error estimates with first-order implicit time-stepping were developed in [undefe, undefa, undef]. Alternatively, goal-oriented adaptivity always requires the solution of an adjoint (dual) problem [undefd, undefc]. Recently, various a posteriori error estimates were derived for semi-discrete formulations with anisotropic mesh refinement either in the $L^{2}(0,T;H^{1}(\Omega)$-norm [undefah, undefu] or in a “damped energy norm” [undefi]; the latter recently led to fully discrete error estimates [undefj] yet under a fixed mesh assumption. Unconditionally stable space-time Galerkin discretization of the second-order wave equation were also derived in [undefr] and, for star-shaped spatial domains, in [undeff] . In [undefao], an adaptive algorithm was proposed for a boundary-element based space-time discretization in one space dimension without mesh-change. Recently, a space-time adaptive algorithm was proposed for the wave equation using an implicit time-stepping strategy [undefo].
@@
In the design of space-time adaptive methods, a posteriori error estimates of fully discrete formulations in time-dependent problems need to include the added effect on the error due to mesh change from one time-step to the next; in fact, some of the above cited works address that particular issue for parabolic problems – see also [undefp, undefae, undefaj]. For both accuracy and efficiency, it is indeed expected, often even required, for an adaptive method to locally refine and coarsen the mesh repeatedly during the entire simulation. Although quantification of mesh-change error in second order hyperbolic problems is less studied, a notable result in this direction was provided by [undefac] in an a priori setting.
@@
To overcome the above mentioned difficulties, we recently derived a posteriori error estimates for the wave equation with explicit LF time integration and mesh change [undefv], starting from the a posteriori analysis for semi-discrete formulations [undeft] (continous in space) discretized in time using a two-step Newmark (or cosine) family of methods, which include the explicit leapfrog method. Moreover, our estimates also allow for LF based explicit local time-stepping (LF-LTS), which circumvents the local CFL bottleneck by taking smaller local time-steps, but only where needed, without sacrificing explicitness.
@@
Here we propose an adaptive strategy for the solution of the time-dependent wave equation based on rigorous a posteriori error estimates which include both mesh change and LF based fully explicit local time-stepping. In Section 2, we present the finite element discretization including the Galerkin FE formulation with explicit LF time integration and mesh change. In Section 3, starting from the a posteriori error estimators from [undefv], which we briefly recall in 3.1, we develop a new adaptive algorithm for the numerical solution of the wave equation which combines three essential ingredients: a posteriori estimates for guaranteed error bounds, mesh change to adapt the mesh from one time-step to another, and local time-stepping to retain the efficiency of explicit time integration even in the presence of local mesh refinement. In fact, local error indicators not only permit one to estimate but also to control the effect of mesh change and adaptivity and thus to decide “on the fly” whether to accept or recompute the numerical solution at the current time-step. By restricting mesh change to compatible meshes, as detailed in Section 3.2, we ensure that local refinement never increases the numerical error while keeping under control the inherent but small error from mesh coarsening. Finally, in Section 4, we present four numerical experiments in one and two space dimensions , which confirm the expected convergence order and illustrate the usefulness of our adaptive strategy.
@@ 2 Galerkin FE formulation with explicit time integration and mesh change
For $u_{0}\in H^{1}_{\Gamma_{0}}(\Omega)$, $v_{0}\in L^{2}(\Omega)$ and $f\in L^{2}(0,T;L^{2}(\Omega))$, it is well-known that the wave equation (2.1) in fact has a unique solution $u\in C^{0}(0,T;H^{1}_{\Gamma_{0}}(\Omega))\cap C^{1}(0,T;L^{2}(\Omega))$ [undefaf].
@@
$h^{n}(x):=\operatorname{diam}\bigcap_{\boldsymbol{x}\in K\in\mathcal{M}_{n}}\bar{K}.$ (2.6)
For each $E\in\mathcal{M}_{n}\cup\operatorname{Sides}\mathcal{M}_{n}$ we denote the constant value of $h^{n}$ by $h_{E}$ (which thus depends implicitly on $n$).
With adaptvity in mind, we allow for the case where $\mathcal{M}_{n}$ (and thus $\mathbbmss{V}_{n}$) changes with time under mesh compatibility conditions, which implies that at each point of the domain either $\mathcal{M}_{n-1}$ is a (compatible) refinement of $\mathcal{M}_{n}$ or conversely – see Section 3.2 and [undefad] for details.
@@
To transfer functions between FE spaces during mesh change, for instance, we introduce the FE transfer operator $\Pi_{n}:C^{0}(\Omega)\rightarrow\mathbbmss{V}_{n}$, which may either denote the $L^{2}$-projection or standard Lagrangian interpolation
@@
thanks to Riesz’s representation theorem via
@@
The particular instance of $\widetilde{A_{n}}$ in (2.16) corresponds to the simplest situation with two local time-steps of size $\Delta t/2$ each for each global time-step of size $\Delta t$. By letting $\widetilde{A_{n}}$ denote a generic perturbed bilinear form induced by local time-stepping, our analysis inherently encompasses situations with different coarse-to-fine time-step ratios, too, which may even change from one locally refined subregion to another across a single mesh. In fact, it even includes a hierarchy of locally refined regions, each associated with its own local time-step [undefn]
@@
$U^{0}$ $=\Pi_{0}u_{0}$ (2.17)
$U^{1}$ $=\Pi_{1}\left[U^{0}+\Pi_{0}v_{0}\,\Delta t+\left(F^{0}-\widetilde{A_{0}}U^{0}\right)\Delta t^{2}\right]$
$U^{n+1}$ $=\Pi_{n+1}\left[2\,U^{n}-\Pi_{n}U^{n-1}+\left(F^{n}-\widetilde{A_{n}}U^{n}\right)\Delta t^{2}\right],\>n=1,\ldots,N.$
@@ 3 Space-time adaptive finite element method
We shall now develop a fully adaptive method for the numerical solution of the wave equation with combines three essential ingredients: a posteriori estimates for guaranteed error bounds, mesh change to adapt ”on the fly” the mesh from one time-step to another, and local time-stepping to retain the efficiency of explicit time integration even in the presence of local mesh refinement.
@@ 3.1 A posteriori error estimates
In [undefv] we derived the following a posteriori error estimates for the discretized time-dependent wave equation (2.17) with mesh change:
@@
$\max_{1\leq n\leq N}\|V^{n-\frac{1}{2}}-v(.,t_{n-\frac{1}{2}})\|_{L^{2}(\Omega)}\\ \leq\|\boldsymbol{e}(0)\|_{\text{erg},\mathcal{A}}+C\left\{2\sum_{m=1}^{2N}\zeta^{m}+\max_{1\leq n\leq N}\varepsilon_{1}^{n}\right\},$ (3.2)
where the initial error, $\boldsymbol{e}(0)$, is defined as $\boldsymbol{e}(0):=[u(0)-U^{0},v(t_{\frac{1}{2}})-V^{\frac{1}{2}}]^{\top}$ , and is a computable quantity. Here the potential energy norm is given by
@@
$\langle\phi,\chi\rangle_{\mathrm{erg},\mathcal{A}}:=\langle\mathcal{A}\phi_{0}\mid\chi_{0}\rangle+\langle\phi_{1},\chi_{1}\rangle\text{ for }\phi=\begin{bmatrix}\phi_{0}\\ \phi_{1}\end{bmatrix},\,\chi=\begin{bmatrix}\chi_{0}\\ \chi_{1}\end{bmatrix}\in H^{1}_{\Gamma_{0}}\times L^{2}(\Omega).$ (3.5)
The upper bounds in (3.1) and (3.2) involve standard a posteriori error estimators for elliptic problems with respect to the energy or the $L^{2}$ norm [undefan, undefb] together with additional terms due to time integration or mesh change:
@@
$\vartheta_{1}^{n}(t)$ $:=\Delta t^{2}\begin{cases}\left\|\frac{1}{2}\partial^{2}U^{n}\ell_{n}(t)-\partial V^{n-\frac{1}{2}}q_{n-\frac{1}{2}}(t)\right\|_{L^{2}(\Omega)},\qquad t\in I^{\prime}_{n},\\[10.00002pt] \left\|\frac{1}{2}\partial^{2}U^{n}\ell_{n}(t)-\partial V^{n-\frac{1}{2}}q_{n+\frac{1}{2}}(t)\right\|_{L^{2}(\Omega)},\qquad t\in I^{\prime}_{n+\frac{1}{2}},\end{cases}$ (3.13)
@@ 3.2 Compatible mesh change
Mesh change during any time-dependent simulation cannot be arbitrary without risking a significant loss in accuracy [undefp]. Indeed both mesh coarsening or refinement will generally increase the numerical error when transferring the FE solution via interpolation or projection from one mesh to another. To ensure that local mesh refinement never leads to a loss in accuracy, we restrict mesh change to compatible meshes only. Moreover, we shall allow coarsening only where appropriate to minimize the resulting inherent information loss.
@@
$\beta_{K}^{0}$ $:=$ $\left\|\left[\Pi_{n+1}-\operatorname{Id}\right]y\right\|_{\mathcal{A},K}=\left\|(y-\hat{y})|_{K}\right\|_{\mathcal{A}}$ (3.18)
$\beta_{K}^{1}$ $:=$ $\left\|\left[\Pi_{n+1}-\operatorname{Id}\right]y\right\|_{L^{2}(K)}=\left\|(y-\hat{y})|_{K}\right\|_{L^{2}}.$ (3.19)
@@
Following [undefae] we now exemplify this point by considering the simplest situation of a one-dimensional FE approximation with continuous piecewise linear polynomials. Again let $K_{-}=(x_{m-1},x_{m})$, $K_{+}=(x_{m},x_{m+1})$ both marked for coarsening and thus possibly replaced by the single element $\widehat{K}=(x_{m-1},x_{m+1})$. For $x\in\widehat{K}$ we have
@@
$y(x)-\hat{y}(x)=\left[y_{m}-\frac{1}{2}(y_{m-1}+y_{m+1})\right]\Phi_{m}^{n}(x),\qquad x\in\widehat{K}$, (3.20)
@@
$\beta_{K_{-}}^{1}$ $=$ $\left|y_{m}-\frac{1}{2}(y_{m-1}+y_{m+1})\right|\left\|\Phi^{n}_{m}\right\|_{L^{2}{(K_{-})}}$
$\beta_{K_{+}}^{1}$ $=$ $\left|y_{m}-\frac{1}{2}(y_{m-1}+y_{m+1})\right|\left\|\Phi^{n}_{m}\right\|_{L^{2}{(K_{+})}},$
@@
When the pre-indicators of two neighboring elements $K_{-}$ and $K_{+}$ are sufficiently small, those elements are subsequently replaced by a single coarser element $\widehat{K}$. If standard polynomial interpolation is used for $\Pi_{n+1}$ in the newly coarsened FE space ${\mathbbmss{V}}_{n+1}$, its remaining nodal values will not change. By allowing for coarsening only in those elements where the pre-indicators are small, we thus also avoid any detrimental drastic increase in the mesh-change indicators $\mu_{0}^{n},\mu_{1}^{n}$ and $\mu_{2}^{n}$ from Section 3.1 .